OpenStudy (anonymous):

Given A(2, 3), B(8, 7), and C(6, 1), which coordinate will make AB perpendicular to cd. please show me the steps

hero (hero):

Where's point D?

OpenStudy (anonymous):

it's not in the question

hero (hero):

That doesn't make sense

hero (hero):

If point D doesn't exist, then segment CD doesn't exist.

OpenStudy (anonymous):

YI amTRYING TO FIND the missing coordinate which is D

hero (hero):

Point D could be any such coordinate that would make AB perpendicular to CD. There's more than one coordinate that would make it true.

OpenStudy (eyust707):

1. Find the slope between AB 2. Take the reciprocal and times it by -1. That's the slope of a perpendicular line. 3. Plug point C and the slope you just found into the point slope form of a line. 4. Any points on that line will satisfy the given conditions. Are you given multiple choices?

hero (hero):

Is point D on the segment AB? That would narrow it down a bit.

OpenStudy (eyust707):

yea quite a bit..

OpenStudy (eyust707):

lol

OpenStudy (anonymous):

yes A.D(3,3) B.D(4,4) c.D(8,4) D.D(9,4)

OpenStudy (eyust707):

The slope between AB should be the opposite reciprocal of the slope between CD. Do you know how to find the slope between two points?

OpenStudy (anonymous):

m=2/3

OpenStudy (anonymous):

perpendiclar slope -3/2

OpenStudy (eyust707):

correct

OpenStudy (eyust707):

which point D makes AD have that slope?

OpenStudy (eyust707):

CD*

hero (hero):

Much easier and quicker with graph paper

OpenStudy (eyust707):

eh.... i dunno i can do all that in my head

OpenStudy (anonymous):

slope=-3/2 point=(6,1)

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

is it right

OpenStudy (eyust707):

\[{ 4 -1 \over 4- 6} = {3 \over -2} = {-3 \over 2}\]

OpenStudy (eyust707):

no need to use the line equation all though what you've written is correct.

OpenStudy (anonymous):

how did you get the last step

OpenStudy (eyust707):

I just tried each one... Try A is the slope between (3,3) and (6,1) equal to -3/2?

OpenStudy (eyust707):

1-3 = -2 6-3 = 3 Slope is -2/3

OpenStudy (eyust707):

try next one..

OpenStudy (eyust707):

1-4 = -3 6-4 = 2 slope is -3/2

OpenStudy (eyust707):

it works!

OpenStudy (anonymous):

yes B is the answer right

OpenStudy (anonymous):

Is there any other method to go straight

OpenStudy (anonymous):

r u there

OpenStudy (anonymous):

Thank you

hero (hero):

Imagine if you had graph paper. You could really see directly how AB perp. to CD with no guesswork or pondering involved.

OpenStudy (anonymous):

So I can do this problem on graph paper

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Refer to the attached plot.